Find the area of a parallelogram ABCD given AB=12cm, AD=20cm, and angle A=30 degrees.
Thanks!

Answer by ikleyn(52786)

.
The area of a parallelogram is the product of the lengths of any two adjacent sides and the sine of the angle between them.


    A = 12*20*sin(30°) = 12*20*0.5 = 120 cm^2.